# Automated Quantum Algorithm Synthesis

### v2

@@ IV.2 Search strategy
Evolutionary search is a gradient-free optimisation technique inspired by the process of natural selection. It remains a popular optimisation method due to its flexibility, conceptual simplicity, and competitive results White et al. (2023). Beginning with a collection of randomly generated candidate algorithms, the fittest candidates are selected and used to produce the next generation, growing the population over time. New algorithms are produced by combining two algorithms or by changing a property of a single algorithm. These operations are referred to as mutations. The pseudo-code that summarises our optimisation procedure is given in Algorithm IV.2.
@@
Evolutionary algorithm: An iterative optimization process that maintains a population of candidate algorithms, refining them through selection, mutation, and evaluation. The search continues until a predefined stopping condition is met. Arrows indicate assignment and plus operations indicate appendage.
@@
In the selection step, we choose two candidate algorithms to mutate, line 7 of Alg. IV.2. Here we balance exploration and exploitation, that is trying new algorithms versus further optimising ones that already have a high fitness score compared to the rest of the population. This is done by randomly deciding between two selection methods, namely random selection or tournament selection (i.e. selection according to fitness ranking) which emphasise exploration and exploitation, respectively. Tournament selection depends on selection pressure which determines the percentage of the population considered. A higher selection pressure corresponds to greater exploitation since each candidate competes against a larger subset of the total population. The probability of performing random or tournament selection and the selection pressure are hyper-parameters of the search strategy. These can be changed over the course of an experiment. We begin in a highly exploratory phase followed by an exploitative phase, with the transition between these phases scheduled according to a sine function. By using a periodic function we can alternate between phases of exploration and exploitation during long experiments.
{minted} [highlightlines=3-4]python algorithm = ( QPivot(H, ”1*”) + QPivot(H, ”*1”) # - + QPivot(H, ”1*”) # + + QCycle(X) + QPivot(MCX, ”*1”) + QPivot(MCX, ”*1”) + QPivot(MCX, ”*1”) + Oracle + QCycle(H) ) * r
@@
{minted} [highlightlines=5]python algorithm = ( Oracle + QCycle(H) + QCycle(X) + QPivot(H, ”*1”) # + + QPivot(MCX, ”*1”) + QCycle(x) ) * r
@@
{minted} [highlightlines=10]python algorithm = ( Oracle + QCycle(H) + QCycle(x) + QPivot(H, ”*1”) + QPivot(MCX, ”*1”) + QPivot(MCX, ”*1”) + QPivot(MCX, ”*1”) + QCycle(H) + Oracle # - + Oracle + Oracle + QCycle(H) + QPivot(H, ”*1”) + QCycle(x) + QCycle(H) ) * r
@@
{minted} [highlightlines=2-7]python algorithm = ( # head Oracle + QCycle(H) + QCycle(x) + QPivot(H, ”*1”) + QPivot(MCX, ”*1”) # tail + QPivot(H, ”*1”) + QCycle(x) + QCycle(H) ) * r
@@
The composability of motifs makes it intuitive to construct methods for generating new algorithms from a pair of parent algorithms. In particular, we used a ‘chop-and-join’ mutation which separates each of a pair of algorithms into two pieces, where the division is made at a random point in the sequence of motifs. The first pieces of each algorithm are exchanged to produce two new algorithms. In addition, three types of mutations were used that operate on a single algorithm. The ‘property’ mutation selects a motif at random from the composition of motifs and either modifies a randomly chosen property or replaces the selected motif with a randomly generated one. The ‘dropout’ and ‘insert’ mutations work by removing and inserting a randomly generated motif, respectively. The probability of using a mask, unmask, cycle, or pivot when randomly generating a motif is an another hyper-parameter of the search. If the task uses an oracle then it is treated as a motif and is also assigned a probability of being used. Fig. 5 shows an example of each of these four mutations. Each example corresponds to a mutation that occurred during the experiment discussed in Fig. 3. If the chop-and-join mutation was omitted then such an evolutionary tree could only be a linear sequence of algorithms.
@@ IV.3 Performance estimation
The first contribution to our fitness function is constructed using state overlap. The training set, $T=\cup_{q\in Q}\{(x^{(q)}_{i},y^{(q)}_{i})\}_{i\in I}$, is a collection of input-target pairs, $(x^{(q)}_{i},y^{(q)}_{i})$, which are grouped by $q\in Q$, where $Q$ is a set of circuit sizes. We define the fitness function as follows
$f(m):=\min_{q\in Q}f_{q}(m):=\min_{q\in Q}\left\{\sum_{i\in I}|\braket{y^{(q)}_{i}}{U^{(q)}_{(m)}}{x^{(q)}_{i}}|^{2}\right\}~{},$ (1)
where $U^{(q)}_{(m)}$ is the quantum circuit (unitary operation) associated with the candidate algorithm $m$ instantiated for the relevant circuit size implied by $q$. Here we take the minimum fitness across the set of circuit sizes, but taking the mean is also valid. It turns out that it is sufficient to evaluate an algorithm on a small number of circuit sizes for the general structure to be learnt. In particular, circuits of size $2$, $3$ and $4$ qubits were sufficient for the Deutsch-Jozsa algorithm and QFT, and circuits of size $3$, $4$ and $5$ qubits were sufficient for Grover’s search.
The fitness can be improved by adding penalties to $f_{q}(m)$ that depend on the complexity of $U^{(q)}_{(m)}$. To favour smaller simpler structures, a penalty is added in terms of the number of unitary gates in the circuit $U^{(m)}_{i}$. A penalty is also added for the number of oracle calls and the number of gate parameters, where applicable. Lastly, we add a penalty based on the Jensen–Shannon divergence Lin (1991) which penalises circuits that perform non-uniformly across the elements of the training set. The relative weights of these penalties are a hyper-parameter of the search.
When an algorithm contains parametrised gates, these are trained numerically using gradient decent  Paszke et al. (2017); Bergholm et al. (2018) to maximise the fitness of the candidate. In order to speed up evaluation, we allow parameter training to run for only a few seconds, which further penalises highly parametrised circuits. Parameters are later trained more carefully during post-processing.
@@ IV.4 Tasks
In our approach, setting up the problem is the primary task of the user, while the task of designing the algorithm is delegated to the computer. As we use a bottom-up approach, the user needs to carefully construct training examples that capture the desired behaviour of the algorithm. If required, the user should also define an appropriate oracle. Here we simply used the known oracles. However, many formulations of the problem may need to be tested to arrive at the final algorithm, which is straightforward to do using our method. Below we discuss the problem setup for each task.
The quantum Fourier transform (QFT) Shor (1999) maps an input state to an output state whose amplitudes are related by the discrete Fourier transform. The training examples are constructed using the computational basis states and their corresponding Fourier-transformed counterparts. An additional training pair is added where the input state is given by an equal superposition of all computational basis states. This penalises erroneous relative phase shifts between all other output states.
@@
Grover’s search Grover (1997) provides an optimal method for searching an unstructured database. The oracle marks the state associated with the sought database entry with a phase, namely $\exp(i\pi)=-1$. A training pair is constructed for each of the possible sought entries (computational basis states) for the given problem size $2^{q}$, where $q$ is the number of qubits. A fixed input state of (|$(\frac{\ket{0}+\ket{1}}{\sqrt{2}})^{\otimes q}$ is used, which encodes the assumption that each database entry has an equal probability being the sought entry.
Oracles are implemented as Python lambda functions so that relevant parameters can be passed at the time of evaluation. Namely the target state, in the case of Grover’s search algorithm, and the function $f$ in the case of the Deutsch-Jozsa algorithm.

### v3

@@ IV.2 Search strategy
Evolutionary search is a gradient-free optimisation technique inspired by the process of natural selection. It remains a popular optimisation method due to its flexibility, conceptual simplicity, and competitive results white2023neural. Beginning with a collection of randomly generated candidate algorithms, the fittest candidates are selected and used to produce the next generation, growing the population over time. New algorithms are produced by combining two algorithms or by changing a property of a single algorithm. These operations are referred to as mutations. The pseudo-code that summarises our optimisation procedure is given in Algorithm IV.2.
@@
Evolutionary algorithm: An iterative optimisation process that maintains a population of candidate algorithms, refining them through selection, mutation, and evaluation. The search continues until a predefined stopping condition is met. Arrows indicate assignment and plus operations indicate appendage.
@@
In the selection step, we choose two candidate algorithms to mutate, line 7 of Alg. IV.2. Here we balance exploration and exploitation, that is trying new algorithms versus further optimising ones that already have a high fitness score compared to the rest of the population. This is done by randomly deciding between two selection methods, namely random selection or tournament selection (i.e. selection according to fitness ranking) which emphasise exploration and exploitation, respectively. Tournament selection depends on selection pressure which determines the percentage of the population considered. A higher selection pressure corresponds to greater exploitation since each candidate competes against a larger subset of the total population. The probability of performing random or tournament selection and the selection pressure are hyper-parameters of the search strategy. These can be changed over the course of an experiment. We begin in a highly exploratory phase followed by an exploitative phase, with the transition between these phases scheduled according to a sine function. By using a periodic function, we can alternate between phases of exploration and exploitation during long experiments.
algorithm = ( QPivot(H, ”1*”) + QPivot(H, ”*1”) # - + QPivot(H, ”1*”) # + + QCycle(X) + QPivot(MCX, ”*1”) + QPivot(MCX, ”*1”) + QPivot(MCX, ”*1”) + Oracle + QCycle(H) ) * r
@@
algorithm = ( Oracle + QCycle(H) + QCycle(X) + QPivot(H, ”*1”) # + + QPivot(MCX, ”*1”) + QCycle(x) ) * r
@@
algorithm = ( Oracle + QCycle(H) + QCycle(x) + QPivot(H, ”*1”) + QPivot(MCX, ”*1”) + QPivot(MCX, ”*1”) + QPivot(MCX, ”*1”) + QCycle(H) + Oracle # - + Oracle + Oracle + QCycle(H) + QPivot(H, ”*1”) + QCycle(x) + QCycle(H) ) * r
@@
algorithm = ( # head Oracle + QCycle(H) + QCycle(x) + QPivot(H, ”*1”) + QPivot(MCX, ”*1”) # tail + QPivot(H, ”*1”) + QCycle(x) + QCycle(H) ) * r
@@
The composability of motifs makes it intuitive to construct methods for generating new algorithms from a pair of parent algorithms. In particular, we used a ‘chop-and-join’ mutation which separates each of a pair of algorithms into two pieces, where the division is made at a random point in the sequence of motifs. The first pieces of each algorithm are exchanged to produce two new algorithms. In addition, three types of mutations were used that operate on a single algorithm. The ‘property’ mutation selects a motif at random from the composition of motifs and either modifies a randomly chosen property or replaces the selected motif with a randomly generated one. The ‘dropout’ and ‘insert’ mutations work by removing and inserting a randomly generated motif, respectively. The probability of using a mask, unmask, cycle, or pivot when randomly generating a motif is another hyper-parameter of the search. If the task uses an oracle then it is treated as a motif and is also assigned a probability of being used. Fig. 5 shows an example of each of these four mutations. Each example corresponds to a mutation that occurred during the experiment discussed in Fig. 3. If the chop-and-join mutation was omitted then such an evolutionary tree could only be a linear sequence of algorithms.
@@ IV.3 Performance estimation
The first contribution to our fitness function is constructed using state overlap. The training set, $T=\cup_{q\in{\mathcal{N}}}\{(x^{(q)}_{i},y^{(q)}_{i})\}_{i\in I}$, is a collection of input-target pairs, $(x^{(q)}_{i},y^{(q)}_{i})$, which are grouped by $q\in{\mathcal{N}}$, where $\mathcal{N}$} is a set of circuit sizes. We define the fitness function as follows
$f({\mathcal{A}})$ $:=\min_{q\in{\mathcal{N}}}f_{q}({\mathcal{A}})$ (2)
$:=\min_{q\in{\mathcal{N}}}\left\{\sum_{i\in I}|\braket{y^{(q)}_{i}|{U([1,\dots,q],\mathcal{A})}|x^{(q)}_{i}}|^{2}\right\}\,$, (3)
where $U([1,\dots,q],\mathcal{A})$ (Eq. (1)) is the quantum circuit (unitary operation) associated with the candidate algorithm $\mathcal{A}$} instantiated for the relevant circuit size implied by $q$. Here, we take the minimum fitness across the set of circuit sizes, but taking the mean is also valid. It turns out that it is sufficient to evaluate an algorithm on a small number of circuit sizes for the general structure to be learnt. In particular, circuits of size $2$, 33 and $4$ qubits were sufficient for the Deutsch-Jozsa algorithm and QFT, and circuits of size $3$, 44 and $5$ qubits were sufficient for Grover’s search.
The fitness can be improved by adding penalties to $f_{q}({\mathcal{A}})$ that depend on the complexity of $U([1,\dots,q],\mathcal{A})$}. To favour smaller simpler structures, a penalty is added in terms of the number of unitary gates in the circuit $U([1,\dots,q],\mathcal{A})$}. A penalty is also added for the number of oracle calls and the number of gate parameters, where applicable. Lastly, we add a penalty based on the Jensen–Shannon divergence lin1991divergence which penalises circuits that perform non-uniformly across the elements of the training set. The relative weights of these penalties are a hyper-parameter of the search.
When an algorithm contains parametrised gates, these are trained numerically using gradient descent  paszke2017automatic; bergholm2018pennylane to maximise the fitness of the candidate. In order to speed up evaluation, we allow parameter training to run for only a few seconds, which further penalises highly parameterised circuits. Parameters are later trained more carefully during post-processing.
@@ IV.4 Tasks
In our approach, setting up the problem is the primary task of the user, while the task of designing the algorithm is delegated to the computer. As we use a bottom-up approach, the user needs to carefully construct training examples that capture the desired behaviour of the algorithm. If required, the user should also define an appropriate oracle. Here, we simply used the known oracles. However, many formulations of the problem may need to be tested to arrive at the final algorithm, which is straightforward to do using our method. Below, we discuss the problem setup for each task.
The quantum Fourier transform (QFT) shor1999polynomial maps an input state to an output state whose amplitudes are related by the discrete Fourier transform. The training examples are constructed using the computational basis states and their corresponding Fourier-transformed counterparts. An additional training pair is added where the input state is given by an equal superposition of all computational basis states. This penalises erroneous relative phase shifts between all other output states.
@@
[…]
Oracles are implemented as Python lambda functions so that relevant parameters can be passed at the time of evaluation. Namely, the target state, in the case of Grover’s search algorithm, and the function $f$ in the case of the Deutsch-Jozsa algorithm.
@@ V Data Availability
VI Code Availability
The underlying code for this study can be accessed at https://github.com/AmyRouillard/qalgosynth.
VII Acknowledgements
The computations were performed at University of Geneva using Baobab HPC service. AR acknowledges the financial assistance of the National Research Foundation (NRF) of South Africa. Opinions expressed and conclusions arrived at, are those of the author and are not necessarily to be attributed to the NRF. ML acknowledges the financial support of the Oppenheimer Memorial Trust of South Africa. ML also acknowledges the financial support of the Unitary Foundation for the development of the Hierarqcal Python package. The funder played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript.
VIII Author Contributions
AR and ML conceived and designed the study. AR implemented the search strategy, performed the numerical experiments and analysed the results. AR prepared the manuscript with input from ML. FP supervised the research. All authors read and approved the final manuscript.
IX Competing interests
All authors declare no financial or non-financial competing interests.
Appendix A Introduction to the DSL
A motif $M_{m}$ is a building block of the DSL and is used to generate a hypergraph $\mathcal{G}_{m}=(Q,E_{m})$, associated with a unitary operation $U^{(m)}$. A circuit diagram can be interpreted as a visual representation of a hypergraph called a parallel aggregated ordered hypergraph (PAOH) valdivia_analyzing_2021. PAOH represents nodes as parallel horizontal lines, and edges as vertical lines, using dots to depict the connections to one or more nodes. Comparing this to a quantum circuit diagram, we can associate qubits with nodes and unitary gates with edges. Using the typical visual conventions for depicting quantum gates provides additional information but does not change the interpretation of the circuit as a PAOH.
Each motif type corresponds to a different method of edge generation. Motifs can be composed to create a sequence of motifs, which can in turn be composed to form higher-level motifs, and so on. Composition is discussed in Sec. A.3 and forms the basis for building candidate algorithms $\mathcal{A}$, which are a sequence of $p$ motifs, $(M_{1},M_{2},\dots,M_{p})$. For the node (qubit) indices $Q$, the quantum circuit is generated according to Eq. (1). The order of gate execution matters for quantum circuits, and gates are added to the circuit from left to right as edges are generated. The importance of edge ordering differentiates the circuit representation from the PAOH. Below, we provide specific examples that illustrate the functionality of each motif in the DSL.
A.1 Cycle
As the name suggests, the cycle, represented as QCycle in the DSL, is a motif that generates a sequence of gates by cycling through the available qubits. Fig. 6 shows the default behaviour of the cycle when used with a $3$-qubit gate. The cycle has a mapping parameter that defines the operation associated with each edge. In this example, the mapping used is the Toffoli gate. A mapping can also be a parametrised gate, such as a parametrised rotation or phase gate. The list of edges generated by the cycle shown in Fig. 6 is $E=[(1,2,3),(2,3,4),\dots,(q-2,q-1,q),(q-1,q,1),(q,1,2)]$, where $q$ is the number of qubits. By default, the cycle uses periodic boundary conditions, so that a unitary gate’s inputs are assigned to qubits modulo $q$. Fig. 7 shows the behaviour of the cycle when these boundary conditions are changed from periodic to open. For open boundary conditions, the list of edges is given by $E=[(1,2,3),(2,3,4),\dots,(q-2,q-1,q)]$.
Qcycle( mapping=Toffoli, stride=1, step=1, offset=0, boundary=”periodic”, edge_order=[1], share_weights = False, )
Figure 6: Default behaviour of the cycle: The DSL motif QCycle— (top) produces the quantum circuit (bottom). The mapping used is the Toffoli gate. The gate inputs and outputs are labelled 0, 11, 22 to illustrate how the ordering is preserved when the cycle of gates wraps around from the bottom to the top of the circuit.
Qcycle( mapping=Toffoli, stride=1, step=1, offset=0, boundary=”open”, edge_order=[1], share_weights = False, )
Figure 7: Cycle with open boundary condition: The DSL motif QCycle— (top) produces the quantum circuit (bottom). Compared to Fig. 6, the gates no longer wrap around from the bottom to the top of the circuit as a result of the open boundary condition.
The stride, step, and offset parameters take on integer values and control how the unitary gate inputs are assigned to qubits, i.e., how edges are generated. The offset controls where the assignment begins within the available qubits. An offset of $1$ effectively shifts all gates down by one qubit, as demonstrated in Fig. 8. Here, the list of edge is given by $E=[(2,3,4),(3,4,5),\dots,(q-2,q-1,q),(q-1,q,1),(q,1,2)]$. Adding an offset does not necessarily preserve the number of gates (edges) generated for two circuits of the same size. For example, Fig. 6 has one more gate than Fig. 8.
Qcycle( mapping=Toffoli, stride=1, step=1, offset=1, boundary=”periodic”, edge_order=[1], share_weights = False, )
Figure 8: Cycle with offset: The DSL motif QCycle— (top) produces the quantum circuit (bottom). The gates are placed with an offset of $1$ qubit with respect to their default position, shown in Fig. 6.
The step parameter, like the offset, controls the traversal of the available qubits. A step of $1$ results in every qubit being visited, if we ignore the effect of the offset. While a step of $n$ results in every $n$th qubit being visited, starting from the first (plus the offset) and proceeding to the last. Fig. 9 illustrates the effect of changing the default step to a value of $2$. In this example, the list of edges is given by $E=[(1,2,3),(3,4,5),\dots,(q-2,q-1,q),(q,1,2)]$.
Qcycle( mapping=Toffoli, stride=1, step=2, offset=0, boundary=”periodic”, edge_order=[1], share_weights = False, )
Figure 9: Changing the step of the cycle: The DSL motif QCycle— (top) produces the quantum circuit (bottom). The step parameter controls which qubits become the first node in each edge. Changing the step from $1$ to $2$ results in a gate being generated at every second qubit. Increasing the step to $3$ would result in a gate at every third qubit, and so on.
The stride controls the interval between node indices connected by the same edge, that is, the intervals between qubit indices assigned to the same unitary gate. For example, a stride of $1$, as in the default case, results in edges being generated for qubit indices that are a distance of $1$ apart. Increasing the stride results in an increased uniform distance between qubit index assignments. Fig. 10 illustrates the gate placement for a stride of $2$ and a step of $1$. In this example, the list of edges is given by $E=[(1,3,5),(2,4,6),\dots,(q-2,q,2),(q-1,1,3),(q,2,4)]$.
Qcycle( mapping=Toffoli, stride=2, step=1, offset=0, boundary=”periodic”, edge_order=[1], share_weights = False, )
Figure 10: Changing the stride of the cycle: The DSL motif QCycle— (top) produces the quantum circuit (bottom). Increasing the stride increases the distance between nodes connected by the same edge.
The edge order parameter controls the order of the gates (edges) in the sequence. The order can be specified using a list of integers or by using one of the two special inputs: [1] indicating default ordering, and [-1] indicating reverse ordering. Fig. 11 demonstrates the reverse ordering functionality, where the list of edges is given by $E=[(q,1,2),(q-1,q,1),(q-2,q-1,q),\dots,(2,3,4),(1,2,3)]$.
Qcycle( mapping=Toffoli, stride=1, step=1, offset=0, boundary=”periodic”, edge_order=[-1], share_weights = False, )
Figure 11: Cycle with reversed edge order: The DSL motif QCycle— (top) produces the quantum circuit (bottom). Compared to Fig. 6, the gates appear in reverse order.
Based on these examples, it is intuitive to understand how gate (edge) generation will change when the unitary mapping used is a one-, two-, or $n$-qubit gate. Notably, changing the stride when a one-qubit gate is used does not affect the gate placement. Each edge connects to a single qubit (node) since a one-qubit gate has a single input. In this case, edge generation is only affected by the offset and step parameters.
Finally, the share weights Boolean parameter sets all gate parameters identically if set to true. This can significantly reduce the number of trainable parameters in the circuit. For non-parametric gates such as the Toffoli gate, changing share weights does not have an effect.
A.2 Pivot
The pivot motif, implemented as Qpivot in the DSL, is so named because it connects a set of pivot qubits to the set of remaining available qubits. Like the cycle, the pivot has a mapping parameter that defines the operation associated with each edge. Pivot qubits are selected using the global pattern parameter. The merge within parameter controls which gate inputs are assigned to qubits in the pivot set, with the remainder being assigned to non-pivot qubits. Both of these parameters make use of bit strings, where a $1$ indicates a pivot qubit.
The DSL is also equipped with the following useful wildcard conventions: a star, *, is filled with zeros, while an exclamation mark, !, is filled with ones. For example, as in Fig. 12, a global pattern of 1* is interpreted as 100...0, so that the motif pivots on the first qubit in the circuit. The remaining gate inputs are assigned to available qubits not in the pivot set, following the same edge generation rules as the cycle. By setting the merge within parameter to 1*, the first gate input is assigned to the pivot qubit(s). Changing the global pattern from 1* to *1 results in a pivot on the last qubit in the circuit, as shown in Fig. 13. In addition, a string not containing * or ! will be assumed to be repeating. For example, 01 is expanded to the required length as 010101..., and truncated from the right if necessary.
Qpivot( mapping=Toffoli, global_pattern=”1*”, merge_within=”1*”, strides=[1, 1, 0], steps=[1, 1, 1], offsets=[0, 0, 0], boundaries=[”open”, open”, periodic”], edge_order=[1], )
Figure 12: Default behaviour of the pivot: The DSL motif Qpivot— (top) produces the quantum circuit (bottom). The global pattern sets the first qubit in the circuit as the pivot qubit. Each gate is placed by cycling through the remaining qubits while making a fixed assignment of the first input of the gate to the top qubit, as specified by the merge within pattern.
The behaviour of the pivot can be specified more precisely as follows. The merge within parameter subdivides a unitary gate’s inputs into two sets: those that will be assigned to pivot qubits and those that will be assigned to the rest. The global pattern similarly subdivides the available qubits (nodes) into pivot and non-pivot qubits. Edges are generated for each of these subsets using the cycle generation pattern. Let us call these the pivot cycle and rest cycle, respectively. Edge generation is completed by merging the two cycles, where this merger occurs using the same logic as used to generate a cycle. For this reason, the pivot requires three values for the strides, steps, offsets, and boundaries parameters. These three values control the behaviour of the pivot cycle, rest cycle, and merger, respectively. During the merger, we step through the edges associated with the rest cycle and join them to edges from the pivot cycle, creating edges that connect the appropriate number of nodes. For example, in Fig. 12, where we pivot on the first qubit, the pivot cycle has edge list $E_{p}=[(1)]$ and the rest cycle an edge list $E_{r}=[(2,3),(3,4),\dots,(q-2,q-1),(q-1,q)]$. These are merged into the final edge list $E=[(1,2,3),(1,3,4),\dots,(1,q-2,q-1),(1,q-1,q)]$. While in Fig. 13, where we pivot on the last qubit, the pivot cycle has edge list $E_{p}=[(q)]$ and the rest cycle an edge list $E_{r}=[(1,2),(2,3),\dots,(q-2,q-1)]$. These are combined into the final edge list $E=[(q,1,2),(q,2,3),\dots,(q,q-2,q-1)]$.
In Fig. 14, two pivot qubits have been specified and the merge within pattern is 1*. Therefore, a rest cycle is generated for an effective two-qubit gate (second and third input of the Toffoli gate) on the subset of non-pivot qubits, with edge list $E_{r}=[(3,4),(4,5),\dots,(q-2,q-1),(q-1,q)]$. The pivot cycle is generated for an effective one-qubit gate (first input of the Toffoli gate) on the two selected pivot qubits, with edge list $E_{p}=[(1),(2)]$. These are merged using periodic boundary conditions to produce the final edge list $E=[(1,3,4),(2,4,5),\dots,(1,q-2,q-1),(2,q-1,q)]$.
Similarly, for a single pivot qubit, we can introduce a step of $2$ in the rest cycle, as shown in Fig. 15 by setting the second element of the steps parameter to $2$. In this example, the pivot cycle has edge list $E_{p}=[(1)]$ and the rest cycle edge list $E_{r}=[(2,3),(4,5),\dots,(q-3,q-2),(q-1,q)]$, with final edge list $E=[(1,2,3),(1,4,5),\dots,(1,q-3,q-2),(1,q-1,q)]$.
To avoid unexpected behaviour, the number of $1$s in the global pattern is systematically increased to match the number of $1$s in merge within, if this is not already the case. For example, a merge within pattern of 111 and a global pattern of 1* are interpreted as 111 and 111*, respectively.
Any single gate placement can be achieved using the pivot motif with an appropriate global pattern. This ensures that it is always possible to map an arbitrary gate-based circuit representation into the DSL. For example, as shown in Fig. 16, the pivot is used to place a single $3$-qubit gate at the second, fourth, and last qubit. The merge within pattern is set to !, which encodes all $1$s. Therefore, all gate inputs are assigned to qubits from the pivot set and none elsewhere. Finally, the edge order parameter changes the order of the gates, identically to the edge order for the cycle. For this reason, an example is not shown.
A.3 Motif composition
The DSL allows motifs to be composed using both the plus and multiplication operations, as shown in Fig. 17. The cycle of Hadamard gates is placed to the left of the cycle of controlled-not gates according to the order of operations in the composition of motifs. This entire structure is repeated three times by multiplying the composition by $3$. We can consider this a first example of a candidate algorithm, and it represents the typical way in which algorithms are constructed in the DSL.
A.4 Mask and Unmask
The Qmask and Qunmask motifs serve to make qubits available and unavailable. It can be useful to make a subset of qubits temporarily unavailable, such as ancillary qubits. The mask and unmask also use a global pattern, which specifies which qubits are to be masked or unmasked, respectively. Fig. 18 shows the behaviour of the mask and can be directly compared to Fig. 17. The unmask has a special global pattern "previous" which undoes the operation of the previous mask, as shown in Fig. 19.
Qpivot( mapping=Toffoli, global_pattern=”*1”, merge_within=”1*”, strides=[1, 1, 0], steps=[1, 1, 1], offsets=[0, 0, 0], boundaries=[”open”, open”, periodic”], edge_order=[1], )
Figure 13: Controlling the pivot qubits using the global pattern: The DSL motif Qpivot— (top) produces the quantum circuit (bottom). Instead of pivoting on the first qubit, as seen in Fig. 13, we now pivot on the last qubit in the circuit. This results in the last qubit being assigned to the first input of each generated gate.
Qpivot( mapping=Toffoli, global_pattern=”11*”, merge_within=”1*”, strides=[1, 1, 0], steps=[1, 1, 1], offsets=[0, 0, 0], boundaries=[”open”, open”, periodic”], edge_order=[1], )
Figure 14: Pivoting on more than one qubit: The DSL motif Qpivot— (top) produces the quantum circuit (bottom). In this example, there are two pivot qubits specified by the global pattern. A first set of edges is generated by a cycle for a one-qubit gate on the two pivot qubits. A second set of edges is generated by a cycle for a two-qubit gate on the remaining available qubits. The final edges are generated by merging edges from these sets, where elements of the second set are merged with elements of the first set cyclically until all elements of the second set are exhausted.
Qpivot( mapping=Toffoli, global_pattern=”1*”, merge_within=”1*”, strides=[1, 1, 0], steps=[1, 2, 1], offsets=[0, 0, 0], boundaries=[”open”, open”, periodic”], edge_order=[1], )
Figure 15: Additional pivot controls: The DSL motif Qpivot— (top) produces the quantum circuit (bottom). The behaviour of the pivot cycle, rest cycle, and merger is controlled by the elements of the steps, strides, offsets, and boundaries parameters. By setting the second element of steps to $2$, the rest cycle now uses a step of $2$, compared to a step of $1$ in Fig. 12.
Qpivot( mapping=Toffoli, global_pattern=”0101*1”, merge_within=”!”, strides=[1, 1, 0], steps=[1, 1, 1], offsets=[0, 0, 0], boundaries=[”open”, open”, periodic”], edge_order=[1], )
Figure 16: Using the pivot to place a single gate: The DSL motif Qpivot— (top) produces the quantum circuit (bottom). By specifying a global pattern containing as many 1’s as the size of the gate, a single gate can be placed on specific qubits. In this case, the three-qubit gate is connected to the second, fourth and last qubit.
[…]
Figure 17: Composing motifs: On the left, two cycles are added together and the resulting motif multiplied by three. On the right is the equivalent circuit representation, in which a cycle of Hadamard gates is followed by a cycle of controlled-not operations, repeated three times. Where motif parameters are unspecified, they are assumed to use default values, as specified in Fig. 6 and Fig. 12.
( Qcycle(mapping=H) + Qcycle(mapping=CNOT) + Qmask( global_pattern=”1*” ) ) * 3
Figure 18: Composing motifs with mask operations: On the left, two cycles and a mask are added together and repeated three times. This results in the circuit on the right, where, with each repetition, the first qubit in the circuit is made unavailable. Where motif parameters are unspecified, they are assumed to use default values, as specified in Fig. 6 and Fig. 12.
( Qcycle(mapping=H) + Qcycle(mapping=CNOT) + Qmask(global_pattern=”1*”) ) * 3 + ( Qunmask( global_pattern=”previous ) ) * 2 + Qcycle(mapping=H)
Figure 19: Unmasking previously masked qubits: Using the global pattern ”previous”—, we can undo the most recent mask operation to make qubits available again. Where motif parameters are unspecified, they are assumed to use default values, as specified in Figures 6 and 12.